Assume that when adults with smartphones are randomly​ selected, 59​% use them in meetings or classes. If 15 adult smartphone users are randomly​ selected, find the probability that exactly 10 of them use their smartphones in meetings or classes.


1
Expert's answer
2021-08-30T17:19:44-0400

Let X will be the number of adults with smartphone which use them in meetings or classes.

X~Binomial(15, 0.59)

"P(X=x) = C^{15}_x(0.59)^x(1-0.59)^{15-x}"

You have to find the probability that exactly 10 of them use their smartphones in meetings or classes.

"P(X=10) = C^{15}_{10} (0.59)^{10} (1-0.59)^{15-10} \\

= \frac{15!}{10!(15-10)!}(0.59)^{10}(0.41)^5 \\

= 3003 \times 0.00511 \times 0.01158 \\

= 0.17782"

Answer: P=0.17782